Ask
your own question, for FREE!
Mathematics
10 Online
A standard deck of playing cards contains 52 cards in 4 suits of 13 cards each. 2 suits are read and 2 suits are black. Find each probability. Assume the first card is replaced before the second card is drawn. P(black, queen)
Still Need Help?
Join the QuestionCove community and study together with friends!
4. P(black, queen) = 26/52 * 4/52 = 0.0385 5. P(black, diamond) = 26/52 * 13/52 = 0.1250 6. P(jack, queen) = 4/52 * 4/52 = 0.0059 hope this helps
thank you very much!
I looked at this and now I just remembered what I thought in my mind when I first came here and you replied all in one time... Here's the question: How did you know I was asking for these answers? Bonus question (I would LOVE you to answer this one: Did and were you sTaLkInG me?
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
17 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal